# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 5.1.1 Checking the $SU(6)_{1}\cong(SU(3)_{2}\times SU(2)_{3})/\mathcal{A}$ Duality
Line label
Quantum Dimension
Conformal Weight
$\mathbf{1}$
$d_{\mathbf{1}}=1$
$h_{\mathbf{1}}=0$
$\mathbf{3}$
$d_{\mathbf{3}}=\frac{1+\sqrt{5}}{2}$
$h_{\mathbf{3}}=4/15$
$\mathbf{\bar{3}}$
$d_{\mathbf{\bar{3}}}=\frac{1+\sqrt{5}}{2}$
$h_{\mathbf{\bar{3}}}=4/15$
$\mathbf{8}$
$d_{\mathbf{8}}=\frac{1+\sqrt{5}}{2}$
$h_{\mathbf{8}}=3/5$
$\mathbf{6}$
$d_{\mathbf{6}}=1$
$h_{\mathbf{6}}=2/3$
$\mathbf{\bar{6}}$
$d_{\mathbf{\bar{6}}}=1$
$h_{\mathbf{\bar{6}}}=2/3$
@@
$SU(6)_{1}$
Line label
Quantum Dimension
Conformal Weight
0
$d_{1}=1$
$h_{1}=0$
1
$d_{2}=1$
$h_{2}=5/12$
2
$d_{3}=1$
$h_{3}=2/3$
3
$d_{4}=1$
$h_{4}=3/4$
4
$d_{5}=1$
$h_{5}=2/3$
5
$d_{6}=1$
$h_{6}=5/12$
@@ 5.2 Further Conformal Embeddings
Notice that the first of these corresponds to the conformal embedding used in [75] in the context of 2D CFT to propose the deep IR of 2D Adjoint QCD. Here we use the conformal embedding to establish a duality of 3D TQFTs. There are various checks that can be performed in the series above. The case $N=4$ in (5.40) gives $Spin(9)_{1}\cong Spin(4)_{6}$, which we verify by a explicit computation in appendix C.2 and is actually equivalent to the duality (5.27) for $N=3$.
@@ 5.2.1 Three-State Potts Model Maverick Coset from Level-Rank Duality
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
2 $\cong$ (4,0)
$d_{2}=(3+\sqrt{5})/2$
$h_{2}=1/5$
$4_{1}$cong (2,4)
$d_{4_{1}}=(1+\sqrt{5})/2$
$h_{4_{1}}=3/5$
$4_{2}$cong (3,3)
$d_{4_{2}}=(1+\sqrt{5})/2$
$h_{4_{2}}=3/5$
@@ A MTC is then a ribbon category, subject to the assumptions stated above, and such that the matrix
with entries on the simple objects is non-degenerate. [^17] Here we should be more precise with the definition of the trace. Intuitively, it is clear that we have to take the loops of the anyons involved. See [22] for a more precise version of this statement. Imposing that this matrix is unitary, up to some conventional overall factor, one obtains the definition of a unitary MTC. In physics one does not quite use $s$, but rather $S=S_{0,0}\,s$, with $S$ unitary and $S_{0,0}=1/\mathcal{D}$, where $\mathcal{D}$ is the total quantum dimension (see below). The modular $S$-matrix satisfies $S_{ab}=S_{ba}=S_{\bar{a}b}^{*}$.
As a special type of monoidal (or fusion) category, a UMTC further consists, in particular, of an associator or $F$-symbols dictating the associativity of line junctions, constrained by the so-called pentagon equations. We will not need these for our applications, so we refer the reader to the references above and the original works [103, 86] for details.
@@
where an anyon and its conjugate satisfy $d_{a}=d_{\bar{a}}$.
@@ B.1 Frobenius Algebras and Local A-Modules
An (associative) [^18] It is known that the theory of local modules can be defined when there is a non-trivial associator defining the algebra (see for example [38]). However, since in (part of) this work we are interested in giving a physical interpretation to the results of [22] in terms of non-invertible symmetries and their relation to TQFT dualities, we restrict to using their definitions and assumptions. algebra (with unit) $\mathcal{A}$ in a Ribbon category $\mathcal{C}$ is a triple $(\mathcal{A},m,\eta)$ consisting of an object $\mathcal{A}$ in $\mathcal{C}$, a multiplication morphism $m\in\mathrm{Hom}(\mathcal{A}\otimes\mathcal{A},\mathcal{A})$ and a unit morphism $\eta\in\mathrm{Hom}(\mathbf{1},\mathcal{A})$ such that
@@
  • A commutative symmetric Frobenius algebra $\mathcal{A}$ has trivial twist $\theta_{\mathcal{A}}=\mathrm{id}_{\mathcal{A}}$. Physically, this mathematical statement captures the fact that only bosons are non-anomalous, and thus, only them can participate in a Frobenius algebra.
@@
These are all the definitions we will need. The main result (or rather, a version of it) for which we need the previous definitions, and upon which the theory of non-invertible anyon condensation rests, is the following rather fundamental theorem of modular tensor categories: If we have a special symmetric commutative Frobenius algebra $\mathcal{A}$ in a modular tensor category $\mathcal{C}$, and if additionally $\mathcal{A}$ is a simple algebra, then the category $\mathcal{C}^{\mathrm{loc}}_{\mathcal{A}}$ of local $\mathcal{A}$-modules is also a modular tensor category. The latter MTC is what we physically refer to as “the gauged theory” or “the condensed theory.”
@@
where the sum runs over all simple objects of $\mathcal{C}$ and $\tilde{c}$ stands for the image of $\mathcal{C}$ in $\bar{\mathcal{C}}$.
@@ C.2 $Spin(9)_{1}\cong\big{(}Spin(3)_{3}\times Spin(3)_{3}\big{)}/\mathcal{A}$
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
1
$d_{1}=\sqrt{2+\sqrt{2}}$
$h_{1}=3/32$
2
$d_{2}=1+\sqrt{2}$
$h_{2}=1/4$
3
$d_{3}=\sqrt{2(2+\sqrt{2})}$
$h_{3}=15/32$
4
$d_{4}=1+\sqrt{2}$
$h_{4}=3/4$
5
$d_{5}=\sqrt{2+\sqrt{2}}$
$h_{5}=35/32$
6
$d_{6}=1$
$h_{6}=3/2$
@@
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
1
$d_{1}=1$
$h_{1}=1/2$
2
$d_{2}=\sqrt{2}$
$h_{2}=9/16$
@@ C.3 $\big{(}SU(4)_{1}\times SU(2)_{-10}\big{)}/\mathcal{A}$
Line label
Quantum Dimension
Conformal Weight
$\mathbf{1}$
$d_{\mathbf{1}}=1$
$h_{\mathbf{1}}=0$
$\mathbf{4}$
$d_{\mathbf{4}}=1$
$h_{\mathbf{4}}=3/8$
$\mathbf{6}$
$d_{\mathbf{6}}=1$
$h_{\mathbf{6}}=1/2$
$\mathbf{\bar{4}}$
$d_{\mathbf{\bar{4}}}=1$
$h_{\mathbf{\bar{4}}}=3/8$
@@
$SU(2)_{10}$
Line label
Quantum Dimension
Conformal Weight
0
$d_{0}=1$
$h_{0}=0$
1
$d_{1}=\sqrt{2+\sqrt{3}}$
$h_{1}=1/16$
2
$d_{2}=1+\sqrt{3}$
$h_{2}=1/6$
3
$d_{3}=\sqrt{2}+\sqrt{2+\sqrt{3}}$
$h_{3}=5/16$
4
$d_{4}=2+\sqrt{3}$
$h_{4}=1/2$
5
$d_{5}=2\sqrt{2+\sqrt{3}}$
$h_{5}=35/48$
6
$d_{6}=2+\sqrt{3}$
$h_{6}=1$
7
$d_{7}=\sqrt{2}+\sqrt{2+\sqrt{3}}$
$h_{7}=21/16$
8
$d_{8}=1+\sqrt{3}$
$h_{8}=5/3$
9
$d_{9}=\sqrt{2+\sqrt{3}}$
$h_{9}=33/16$
10
$d_{10}=1$
$h_{10}=5/2$
Table 15: $SU(2)_{10}$ data.
@@ C.4 $(E_{6})_{1}\cong(G_{2})_{3}/\mathcal{A}$
Line label
Quantum Dimension
Conformal Weight
$\mathbf{1}$
$d_{\mathbf{1}}=1$
$h_{\mathbf{1}}=0$
$\mathbf{7}$
$d_{\mathbf{7}}=\frac{3+\sqrt{21}}{2}$
$h_{\mathbf{7}}=2/7$
$\mathbf{27}$
$d_{\mathbf{27}}=\frac{7+\sqrt{21}}{2}$
$h_{\mathbf{27}}=2/3$
$\mathbf{77}$
$d_{\mathbf{77}}=\frac{3+\sqrt{21}}{2}$
$h_{\mathbf{77}}=8/7$
$\mathbf{14}$
$d_{\mathbf{14}}=\frac{3+\sqrt{21}}{2}$
$h_{\mathbf{14}}=4/7$
$\mathbf{64}$
$d_{\mathbf{64}}=\frac{5+\sqrt{21}}{2}$
$h_{\mathbf{64}}=1$
@@
Line label
Quantum Dimension
Conformal Weight
$\mathbf{1}$
$d_{\mathbf{1}}=1$
$h_{\mathbf{1}}=0$
$\mathbf{27}$
$d_{\mathbf{27}}=1$
$h_{\mathbf{27}}=2/3$
$\overline{\mathbf{27}}$
$d_{\overline{\mathbf{27}}}=1$
$h_{\overline{\mathbf{27}}}=2/3$
@@ C.5 $SU(2)_{3}\cong\big{(}USp(6)_{1}\times SO(3)_{-4}\big{)}/\mathcal{A}$
Line label
Quantum Dimension
Conformal Weight
$\mathbf{1}$
$d_{\mathbf{1}}\ \ =1$
$h_{\mathbf{1}}\ \ =0$
$\mathbf{14}^{\prime}$
$d_{\mathbf{14}^{\prime}}=1$
$h_{\mathbf{14}^{\prime}}=3/4$
$\mathbf{14}$
$d_{\mathbf{14}}\,=\frac{1+\sqrt{5}}{2}$
@@
$\mathbf{6}$
$d_{\mathbf{6}}\ \ =\frac{1+\sqrt{5}}{2}$
$h_{\mathbf{6}}\ \ =7/20$

### v2

@@ 5.1.1 Checking the $SU(6)_{1}\cong(SU(3)_{2}\times SU(2)_{3})/\mathcal{A}$ Duality
Line label Quantum Dimension Conformal Weight
$\mathbf{1}$ $d_{\mathbf{1}}=1$ $h_{\mathbf{1}}=0$
$\mathbf{3}$ $d_{\mathbf{3}}=\frac{1+\sqrt{5}}{2}$ $h_{\mathbf{3}}=4/15$
$\mathbf{\bar{3}}$ $d_{\mathbf{\bar{3}}}=\frac{1+\sqrt{5}}{2}$ $h_{\mathbf{\bar{3}}}=4/15$
$\mathbf{8}$ $d_{\mathbf{8}}=\frac{1+\sqrt{5}}{2}$ $h_{\mathbf{8}}=3/5$
$\mathbf{6}$ $d_{\mathbf{6}}=1$ $h_{\mathbf{6}}=2/3$
$\mathbf{\bar{6}}$ $d_{\mathbf{\bar{6}}}=1$ $h_{\mathbf{\bar{6}}}=2/3$
@@
$SU(6)_{1}$
Line label Quantum Dimension Conformal Weight
0 $d_{1}=1$ $h_{1}=0$
1 $d_{2}=1$ $h_{2}=5/12$
2 $d_{3}=1$ $h_{3}=2/3$
3 $d_{4}=1$ $h_{4}=3/4$
4 $d_{5}=1$ $h_{5}=2/3$
5 $d_{6}=1$ $h_{6}=5/12$
@@ 5.2 Further Conformal Embeddings
Notice that the first of these corresponds to the conformal embedding used in [78] in the context of 2D CFT to propose the IR phases of 2D Adjoint QCD. Here we use the conformal embedding to establish a duality of 3D TQFTs. There are various checks that can be performed in the series above. The case $N=4$ in (5.40) gives $Spin(9)_{1}\cong Spin(4)_{6}$, which we verify by a explicit computation in appendix C.2 and is actually equivalent to the duality (5.27) for $N=3$.
@@ 5.2.1 Three-State Potts Model Maverick Coset from Level-Rank Duality
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
2 $\cong$ (4,0) $d_{2}=(3+\sqrt{5})/2$ $h_{2}=1/5$
$4_{1}$cong (2,4) $d_{4_{1}}=(1+\sqrt{5})/2$ $h_{4_{1}}=3/5$
$4_{2}$cong (3,3) $d_{4_{2}}=(1+\sqrt{5})/2$ $h_{4_{2}}=3/5$
@@ A MTC is then a ribbon category, subject to the assumptions stated above, and such that the matrix
with entries on the simple objects is non-degenerate. [^18] Here we should be more precise with the definition of the trace. Intuitively, it is clear that we have to take the loops of the anyons involved. See [22] for a more precise version of this statement. Imposing that this matrix is non-degenerate, one obtains the definition of a MTC. In physics one does not quite use $s$, but rather $S=S_{0,0}\,s$, with $S$ unitary and $S_{0,0}=1/\mathcal{D}$, where $\mathcal{D}$ is the total quantum dimension (see below). The latter object can be recognized as the standard modular $S$-matrix encountered in physics satisfying $S_{ab}=S_{ba}=S_{\bar{a}b}^{*}$.
As a special type of monoidal (or fusion) category, a UMTC further consists of an associator or $F$-symbols dictating the associativity of line junctions, constrained by the so-called pentagon equations. Unitarity corresponds to the statement (among others) that a basis of the $F$-symbols exists such that the corresponding $F$-matrices are unitary. (See [106] for a more precise discussion of unitarity in the context of ribbon categories.) We will not need all these details for our applications, so we refer the reader to the references above and the original works [107, 89] for details.
@@
where an anyon and its conjugate satisfy $d_{a}=d_{\bar{a}}$. Unitarity requires that quantum dimensions are positive $d_{a}>0$.
@@ B.1 Frobenius Algebras and Local A-Modules
An (associative) algebra (with unit) $\mathcal{A}$ in a Ribbon category $\mathcal{C}$ is a triple $(\mathcal{A},m,\eta)$ consisting of an object $\mathcal{A}$ in $\mathcal{C}$, a multiplication morphism $m\in\mathrm{Hom}(\mathcal{A}\otimes\mathcal{A},\mathcal{A})$ and a unit morphism $\eta\in\mathrm{Hom}(\mathbf{1},\mathcal{A})$ such that
@@
  • A commutative symmetric Frobenius algebra $\mathcal{A}$ has trivial twist $\theta_{\mathcal{A}}=\mathrm{id}_{\mathcal{A}}$. Physically, this mathematical statement captures the fact that only bosons are non-anomalous, and thus, only them can participate in a commutative Frobenius algebra.
@@
These are all the definitions we will need. The main result (or rather, a version of it) for which we need the previous definitions, and upon which the theory of non-invertible anyon condensation rests, is the following rather fundamental theorem of algebras over modular tensor categories: If we have a special symmetric commutative Frobenius algebra $\mathcal{A}$ in a modular tensor category $\mathcal{C}$, and if additionally $\mathcal{A}$ is a simple algebra, then the category $\mathcal{C}^{\mathrm{loc}}_{\mathcal{A}}$ of local $\mathcal{A}$-modules is also a modular tensor category. The latter MTC is what we physically refer to as “the gauged theory” or “the condensed theory.”
@@
where the sum runs over all simple objects of $\mathcal{C}$ and $\tilde{c}$ stands for the image of $\mathcal{C}$ in $\bar{\mathcal{C}}$, where $\bar{\mathcal{C}}$ is the time-reversal of $\mathcal{C}$.
@@ C.2 $Spin(9)_{1}\cong\big{(}Spin(3)_{3}\times Spin(3)_{3}\big{)}/\mathcal{A}$
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
1 $d_{1}=\sqrt{2+\sqrt{2}}$ $h_{1}=3/32$
2 $d_{2}=1+\sqrt{2}$ $h_{2}=1/4$
3 $d_{3}=\sqrt{2(2+\sqrt{2})}$ $h_{3}=15/32$
4 $d_{4}=1+\sqrt{2}$ $h_{4}=3/4$
5 $d_{5}=\sqrt{2+\sqrt{2}}$ $h_{5}=35/32$
6 $d_{6}=1$ $h_{6}=3/2$
@@
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
1 $d_{1}=1$ $h_{1}=1/2$
2 $d_{2}=\sqrt{2}$ $h_{2}=9/16$
@@ C.3 $\big{(}SU(4)_{1}\times SU(2)_{-10}\big{)}/\mathcal{A}$
Line label Quantum Dimension Conformal Weight
$\mathbf{1}$ $d_{\mathbf{1}}=1$ $h_{\mathbf{1}}=0$
$\mathbf{4}$ $d_{\mathbf{4}}=1$ $h_{\mathbf{4}}=3/8$
$\mathbf{6}$ $d_{\mathbf{6}}=1$ $h_{\mathbf{6}}=1/2$
$\mathbf{\bar{4}}$ $d_{\mathbf{\bar{4}}}=1$ $h_{\mathbf{\bar{4}}}=3/8$
@@
$SU(2)_{10}$
Line label Quantum Dimension Conformal Weight
0 $d_{0}=1$ $h_{0}=0$
1 $d_{1}=\sqrt{2+\sqrt{3}}$ $h_{1}=1/16$
2 $d_{2}=1+\sqrt{3}$ $h_{2}=1/6$
3 $d_{3}=\sqrt{2}+\sqrt{2+\sqrt{3}}$ $h_{3}=5/16$
4 $d_{4}=2+\sqrt{3}$ $h_{4}=1/2$
5 $d_{5}=2\sqrt{2+\sqrt{3}}$ $h_{5}=35/48$
6 $d_{6}=2+\sqrt{3}$ $h_{6}=1$
7 $d_{7}=\sqrt{2}+\sqrt{2+\sqrt{3}}$ $h_{7}=21/16$
8 $d_{8}=1+\sqrt{3}$ $h_{8}=5/3$
9 $d_{9}=\sqrt{2+\sqrt{3}}$ $h_{9}=33/16$
10 $d_{10}=1$ $h_{10}=5/2$
Table 15: $SU(2)_{10}$ data.
@@ C.4 $(E_{6})_{1}\cong(G_{2})_{3}/\mathcal{A}$
Line label Quantum Dimension Conformal Weight
$\mathbf{1}$ $d_{\mathbf{1}}=1$ $h_{\mathbf{1}}=0$
$\mathbf{7}$ $d_{\mathbf{7}}=\frac{3+\sqrt{21}}{2}$ $h_{\mathbf{7}}=2/7$
$\mathbf{27}$ $d_{\mathbf{27}}=\frac{7+\sqrt{21}}{2}$ $h_{\mathbf{27}}=2/3$
$\mathbf{77}$ $d_{\mathbf{77}}=\frac{3+\sqrt{21}}{2}$ $h_{\mathbf{77}}=8/7$
$\mathbf{14}$ $d_{\mathbf{14}}=\frac{3+\sqrt{21}}{2}$ $h_{\mathbf{14}}=4/7$
$\mathbf{64}$ $d_{\mathbf{64}}=\frac{5+\sqrt{21}}{2}$ $h_{\mathbf{64}}=1$
@@
Line label Quantum Dimension Conformal Weight
$\mathbf{1}$ $d_{\mathbf{1}}=1$ $h_{\mathbf{1}}=0$
$\mathbf{27}$ $d_{\mathbf{27}}=1$ $h_{\mathbf{27}}=2/3$
$\overline{\mathbf{27}}$ $d_{\overline{\mathbf{27}}}=1$ $h_{\overline{\mathbf{27}}}=2/3$
@@ C.5 $SU(2)_{3}\cong\big{(}USp(6)_{1}\times SO(3)_{-4}\big{)}/\mathcal{A}$
Line label Quantum Dimension Conformal Weight
$\mathbf{1}$ $d_{\mathbf{1}}\ \ =1$ $h_{\mathbf{1}}\ \ =0$
$\mathbf{14}^{\prime}$ $d_{\mathbf{14}^{\prime}}=1$ $h_{\mathbf{14}^{\prime}}=3/4$
@@
$\mathbf{6}$ $d_{\mathbf{6}}\ \ =\frac{1+\sqrt{5}}{2}$ $h_{\mathbf{6}}\ \ =7/20$